OpenStudy (anonymous):

Absolute Value Equations and Inequalities lxl + 5 = 18

OpenStudy (anonymous):

x = 13

OpenStudy (anonymous):

explain?? please.......

OpenStudy (anonymous):

subtract 5 from each side to get x + 5 - 5 = 18 - 5 then simplify to get x = 13

OpenStudy (anonymous):

ok so if u had a problem like this: how would you solve it? lxl + 3 < 5

OpenStudy (anonymous):

Same thing, just be careful you can not have an absolute value equaling a negative number.

OpenStudy (anonymous):

would this be right then? - 8< x <2

OpenStudy (anonymous):

There seems to be a small bit of confusion here. First, every absolute value has 2 solutions. |x|+5 = 18 |x| = 13 At this stage we split the equation into a positive and negative form First, the positive (everything stays the same) x = 13 Second, the negative (multiply the absolute side by -1) -x = 13 x = -13 The 2 answers are x = 13, -13

OpenStudy (anonymous):

The basic solution given is ok, but the absolute value wasn't taken into account |x| = 13, which means x= 13 or =-13

OpenStudy (anonymous):

thank you guys. i appreciate the help :)

OpenStudy (anonymous):

For the second problem |x| + 3 < 5 We simplify it to, |x| < 5 Then we make the same split. First, the positive x < 5 Second, the negative -x < 5 Remember, when you multiply both sides by a negative, the inequality flips, so x > -5 Your answer here then is, -5<x<5

OpenStudy (anonymous):

Oops, I meant simplify to |x| < 2 which then gives -2 < x < 2

OpenStudy (anonymous):

these are my answer choices. - 8< x <2 - 2 < x <8 3< x <5 -3< x <5

OpenStudy (anonymous):

Are you certain you wrote the problem correctly?

OpenStudy (anonymous):

lx +3l <5

OpenStudy (anonymous):

I thought so, the answer in that case is First x + 3 < 5 x < 5 Second -x -3 < 5 -x < 8 x > -8 So, -8 < x < 2.

OpenStudy (anonymous):

Wow, I screwed up that first part again, First, x + 3 < 5 x < 2

OpenStudy (anonymous):

lol thank you mucho!!!